HUBBLE SPACE TELESCOPE DAILY REPORT #3079 PERIOD COVERED: 0000Z (UTC) 03/22/02 - 0000Z (UTC) 03/25/02 Daily Status Report as of 084/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Fourteen Sets of ACS/WFC/HRC 8947 (Weekly Test) The Advanced Camera for Surveys (WFC and HRC) was used to perform basic tests to monitor, the read noise, the development of hot pixels and test for any source of noise in ACS CCD detectors. This program will be executed at least once a day for the entire lifetime of ACS. There were no problems reported. 1.2 Completed Six Sets of STIS/CCD 8964 (CCD Darks) The Space Telescope Imaging Spectrograph (CCD) was used to obtain dark frames twice daily, as in a corresponding calibration program. Standard calibration reference darks are constructed and hot pixel lists are updated for use in pipeline reduction of SMOV3B data. All observations completed without incident. 1.3 Completed Forty-one Sets of NICMOS/1/2/3 8945 (Cooling System Monitoring) The Near Infrared Camera and Multi-Object Camera (NIC1, NIC2 and NIC3 was used 1}.to measure NICMOS detector performance during the cool-down and steady state operation of the NCS. 2}. to demonstrate stability {+/-0.1K} of the NICMOS detector temperature at the optimal science operating temperature. and 3}.to demonstrate repeatability {+/-0.1 K} of NICMOS detector temperature following changes from the optimal science operating temperature. No problems were encountered. 1.4 Completed Six Sets of HST 8968 (Vehicle Disturbance Test) The Vehicle Disturbance Test {VDT} was used to characterize uncompensated environmental disturbances acting upon the HST during normal operation. The VDT is a passive test {not a forced-response test} used to obtain signatures for both externally induced {e.g. SA-3} and internally induced {e.g. NCC and ACS mechanisms} disturbances for comparison with past VDT results. The disturbances observed will be used as the nominal on-orbit disturbances in pointing control simulations until the next VDT is run. The test occurs after release, and most of the VDT can be run during the BEA period. The -V1 sunpoint portion of the VDT occurs after the BEA period is complete. There were no reported problems. 1.5 Completed Two Sets of ACS/WFC/HRC 9013 (Coarse Corrector Alignment) The Advanced Camera for Surveys (WFC and HRC) was used in an attempt to approximately optimize the ACS image quality over the fields of both the WFC and HRC channels by adjusting the IM1 and M1 corrector mechanisms in both tip/tilt {cylinder rotation} and focus. These observations of a moderate density star field constitute the "first light" external images for the HST/ACS. No anomalous activity occurred. 1.6 Completed Six Sets of STIS/CCD 8963 (CCD Bias) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD read noise from multiple bias frames obtained in common science modes: amplifier D, gains 1 and 4 electrons/DN, and binning factors of 1x1, 1x2, 2x1, and 2x2. The read noise measurements before and after NCC turn-on are then compared, and standard calibration reference files for use in pipeline reduction of SMOV3B data are constructed. The proposal had no reported problems. 1.7 Completed Twelve Sets of WF/PC-2 8950 (SM3B Cool Down, Contamination Monitor, and Focus Check) The WF/PC-2 was used to monitor the planned cooldown to -88C, followed by UV monitors, focus checks, as well as decons and associated monitors. No problems were noted. 1.8 Completed Two Sets of FGS/1 9581 (A BEA Science Program: A High Angular Resolution Survey of Some of the Galaxy's Most Massive Stars) Fine Guidance Sensor #1 was used to make optimal use of otherwise unused telescope time while HST's pointing is restricted to the BEA cone during the first 12 days of SMOV3B. The launch date of Feb 29 placed two of the Galaxy's most active star forming regions, Trumplers 14 and 16 in the Carina Nebula, within the BEA for the entire 12 day period. FGS1r was used to conduct a high angular resolution survey of the most massive and luminous stars in these two magnificent clusters to establish the incidence of binarity at the high mass end of the HR diagram. This survey will close the detection gap between spectroscopic and speckle surveys in which duplicity can be concealed and will dramatically improve our knowledge of the IMF at the high mass end of the main sequence as well as the luminosity function {LF} of star forming regions which produce such high mass stars. Finally, this survey will identify targets for an approved SIM Key Science project to dynamically determine O-star masses. There was no anomalous activity. 1.9 Completed WF/PC-2 8951 (SM3B Lyman Alpha Check) The WF/PC-2 was used to perform GRW+70D5824 throughput measurements in UV filters and UV crossed with long-pass filters {to determine red leak contribution} that will be used to indicate any drop in throughput at Lyman alpha, which could be due to contamination in the pick-off mirror {or primary/secondary mirror}. There were no reported problems. 1.10 Completed STIS/CCD 8958 (Image Quality) The Space Telescope Imaging Spectrograph (CCD) was used to check image quality using two tests. In the first test, an ACQ/PEAK was used to dither the 0.1x0.09 aperture across a point source to sample the PSF. Light was dispersed entering the spectrograph onto the CCD using the G230LB grating. Relative counts at different wavelengths will give a measure of PSF shape at different wavelengths. In the second test, a point source is imaged onto the CCD through the F28x50OII filter. The resulting PSF is then compared to the PSF shape prior to SM3B. No problems were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 23 Successful: 23 Scheduled Re-acquisitions: 16 Successful: 15 The re-acquisition at 084/0740Z failed. Two proposals, ACS 8947 and NIC 8945, each occurring during the period of the next report, may have been affected. 2.2 FHST Updates: Scheduled: 45 Successful: 45 2.3 Operations Notes: The NICMOS Cryo cooler cooldown continues nominally. During the period, operations notes were used to: Load TMDIAG Turn Off ACS HRS/SBC Corrector Status Limit Lim Adjust for QDV10GT(1-2) Limit Management During WFPC2 Decontamination Adjust ACS Error Count Limit Turn On ACS HRC/SBC Corrector Status Limit NICMOS MEB Limit Changes Vehicle Disturbance Test #4 executed nominally between 081/034:0Z0 and 081/115000Z. The intent of VDT #4 was to capture HST disturbance characteristics during the post-surging regime of the NCC. Although NCS cooldown is progressing slower than expected, VDT #4 was run anyway. This test was a 5-orbit test that acquired HST dynamic responses due to NCC surging (and other) disturbances at the +73 hour mark since the transition to PID Operate, at the maximum compressor speed during surging. As observed during VDT #3, the fundamental NCC surging disturbance frequency continues to drop as the NCS slowly cools. During VDT #4, the fundamental surging disturbance was observed to be ~1.34 Hz, and integer multiples of the fundamental frequency were measured by the HST rate gyros. Also observed during VDT #4 were similar disturbance events that have been observed in previous disturbance tests, such as reaction wheel zero-speed crossings. Per an operations request, a high/low gyro bias update was performed at 081/2232Z. A TTR was written due to realtime minor frame gaps and data dropouts during a TDE support at 082/0013Z. There was no data loss. ROP IC-2 was used to turn SSA transmitter #1 on for the intervals 081/1251Z to 081/1301Z and 083/1750Z to 083/1802Z. The SSR-1 EDAC error counter was cleared at 081/1302Z per ROP SR-1A. HSTAR 8571 was written when ACS parameters JHFCSSTA and JERRCNT flagged out-of-limits high at 081/150128Z. They returned to normal ranges at 081/151458Z. A status buffer message was received that an HRC corrector focus move was unsuccessful because the final position achieved was not with tolerance of the desired position. The ACS flight software error flag was cleared at 081/2154Z per an operations request. At 082/210208Z, NICMOS parameters NMEBTMP2 (value 30.0645 degC -- limit 30.00) flagged out-of-limits high and at 082/213929Z NRAD2TMP (value 30.0683 degC -- limit 30.00) flagged high. HSTAR 8573 was written. Then, at 083/003000Z, per HSTAR 8574, NMEBTMP2 increased in temperature to 30.5285 degC. Later, at 083/010008Z, NMEBTMP1 flagged out-of-limits high with a maximum reading of 30.0645 degC. For this occurrencHSTAR 8575 was written. Vehicle Disturbance Test #5 executed nominally between 081/224000Z and 082/065000Z. The intent of VDT #5 was to capture HST disturbance characteristics during the post-surging regime of the NCC at V1 sunpoint. Although NCS cooldown is progressing slower than expected, VDT #5 was run anyway. This test was a 5-orbit test that acquired HST dynamic responses due to NCC surging (and other) disturbances at the +92 hour mark since the transition to PID Operate, at the maximum compressor speed during surging (~7100 rps). The structural resonance at ~1.2 Hz observed in the V3-axis during previous VDTs at +V3 sunpoint is now observed in the V1-axis, thereby distinguishing this mode as a Solar Array 3 mode. This mode is probably a solar array in-plane bending mode, a prominent contributor to HST jitter in +V3 sunpoint attitudes. Terminator disturbances at EOD/EON have been observed for the first time during VDT #4 and #5. These disturbances are often not obvious, but several have occurred over the past 24-hours that approach a magnitude of 9 milli-arcseconds (60 second rms). Post-test analysis will be performed to determine when the first of these disturbances occurred, as they were not seen during VDTs 1-3. As observed during VDT #3 and #4, the fundamental NCC surging disturbance frequency continues to drop as the NCS slowly cools. During VDT #5, the fundamental surging disturbance was observed to be ~1.27 Hz, and integer multiples of the fundamental frequency were measured by the HST rate gyros. NCC-induced disturbances during surging are contributing 10% or less to the overall HST jitter, consistent with pre-SM3B analysis predictions. Outside VDT intervals, Fine Guidance Sensors are acquiring guide stars, and HST is performing science observations during NCC surging, again consistent with PCS analytical predictions. The spacecraft attitude used for VDT #5 resulted in high gravity gradient environmental torques about the V2-axis. Several limit violations occurred during VDT #5 for the mnemonic QDV10GT1, the flight computer computation of V2-axis gravity gradient torque. During SM3B, the flight computer database was updated with the new post-SM3B HST inertia tensor, approximately 20% larger than the pre-mission inertia. However, no change was made to the ground system gravity gradient torque limits, which also should have been increased proportionally with the inertia increase. OPS Note #963 was written to set the limits for the maximum V2 and V3-axis gravity gradient torque, and HSTAR 8572 identifies this anomaly. The limit violations had no impact upon VDT #5. Subsequent analysis to resolve this HSTAR will be performed to select the final limit value and assess any impacts the new limits may have upon the HST momentum management system. A TTR was written for a required re-transmit at 084/0319Z during a NSSC-1 load. Accordingly, ROP NS-5 was utilized to reset SI C&DH errors at 084/0404Z. As recorded in HSTAR 8576, the speeds for reaction wheels 1 and 3 broake limits during a slew. The maximum speeds reached was 2300 revolutions per minute. Using ROP DF-18A, the engineering status buffer limit for SESBSLD (=4) was modified at 084/0759Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW